OpenStudy (anonymous):

9x-9y=27 7y-3x=-11 solve using elimination method method

OpenStudy (anonymous):

x=5/2 y=-1/2

OpenStudy (gw2011):

Here is how you solve the problem: Multiply the first equation by 3 and the second equation by 9 and you get: 27x-27y=81 -27x+63y=-99 Add both of the above equations and you get: 36y=-18 y=-18/36=-1/2 Substitute y=-1/2 into one of the equations and you get: -3x +7(-1/2)=-11 -3x-7/2=-11 -3x=-15/2 x=5/2